Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day13


Crew 103 Log Book for April 21, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status: Everybody is fine

Time Departed/Returned from EVA:

16h30-18h45

Brief Narrative of Field Mission Results:

We continued our measurements with the algae's. For the algae experiment with blue LEDs,  we found a more linear thant exponential

growth but we are still waiting for other results.

G. Van Roy, W. de Groot & T. Windels went on EVA to get the missing coordinates.

EVA Data/Interpretations:

The results were given in a previous report. They just have the coordinates now.

Engineering/Hab Maintenance:

nothing to mention

Plans for Tomorrow:

Today was our last day on Mars! So tomorrow we will finish our reports and tidy the Hab!

Support Requested:

/

Miscellaneous:

Mars, we will miss you! ;)



Commander's Report

Anne-Sophie Kalbfleisch Reporting

Today was our last day on Mars... Time goes by so rapidly! Today we

spent a lot of time making reports for our university, filling our psychological reports and we continued our algae measurements. We're really looking forward having all the results!

G. Van Roy, T. Windels and W. De Groot went on EVA to get the coordinates of the places where the soil samples were taken because, as you may remember, the battery of the GPS was flat.

Today was also our last non-cooking day, fortunately because we had

just got enough left for today!

See you tomorrow for the final report :)


Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,46

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,36

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 3 galons approximately

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the all generator off!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 7,5

Hab Potable Water Tank Level (inches from bottom): 7

Potable Water Meter Reading: 45304,0

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: All ventilators are off. All the thermostats are down to 50. One of the propane heaters is still putting out heat. the temperature inside is 81°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no): no

Tire Status: ok

Odometer at end of day: 124250,5

Notes/Comments on Hab Car: /


ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel consumed : 1/4

Tire Status: ok

Comments: no problems with the gearbox anymore


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/4

Tire Status: 2 tires maybe need to be repumped tomorow

Comments: We have to start manually the ATV because the electric button doesn't work anymore.


ATV 3 (Viking I):

Used: yes

Oil Checked: no

Fuel Consumed: 1/4

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 60

Actual temperature upstairs (Farenheit): 76

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 76


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: internet is still running well


General Engineering Notes/Comments: /


Science Report

S.Viroux and T.Windels Reporting

Positioning System Project

As we know, there are no artificial satellites flying around Mars. The utilization of a GPS on Mars is thus impossible. That's why astronauts would need an other system to locate himself.

The purpose of this project is to create a grid composed of many transmitters and one receiver in order to create a triangulation system. the biggest part of our project was the study and the analysis of many electronic components and the choice of a communication protocol (Wifi,Zigbee,…).

The wireless technology chosen to allow the devices to 'talk' to each other is Zigbee technology. Our transmitters and receivers are thus composed of arduino or fio boards and Xbee antennas (we needed an Xbee shield for the arduino board), all battery powered.

After this introduction research, the first step was to succeed in making our transmitters talking to the receiver, and then getting the RSSI (Received Signal Strength Indicator). To achieve this we needed to understand and configure the antennas (change the ID,MY,DL,D0,IT etc…), which took us a few days!

With this RSSI information, we are normally able to determine the distance between the receiver and the transmitter, and then when we take the distance to three transmitters into account, we are also able to locate the receiver (the location of the transmitters is known). In order to deduce the distance from the RSSI signal, we have to use this equation : Pr = Pt + Gt + Gr - 20*log10(lambda/(4*pi*D))^n

With : Pr= received power (RSSI)

      Pt= transmitted power (which is a characteristic of our antenna)

      Gt= transmission profit : 2 dBi

      Gr= received profit : 2dBi

      n : to determine

Processing language was used to implement all those steps.

Unfortunately, for many unknown reasons, we didn't succeed in getting the RSSI signal at the MDRS. The laptop we used in Belgium couldn't be taken here, and even if we had 5 laptops at the MDRS, none of them could run Processing suitably. It's probably an hardware issue but impossible to determine it!

The work that should have been done at the MDRS, would have been to improve the grid and to adjust our formula with several outside tests.


Chef's Report

Gregory Van Roy Reporting

1a) Today was

A non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Nothing special today

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Beerbreak Creamy Potato, Chicken, Potatoes, Hollandaise sauce and

pomegranate drink mix

3) List any main meal foods not finished by the crew (leftovers stored for

later use or discarded)

Potatoes : 3/4 of the pack left

4)  Recipes for anything more complicated than following package

directions exactly, or re-hydrating.

None

5)  Lessons learned  (if any)

None

6) Comments/questions for Kim and Jean, the food study investigators

It was the last day on Mars today so we will take a big steak tomorrow

with lots of fresh stuffs :) But the challenge was really interesting!

Some global comments about our cooking experiment during two weeks :

-Need more pasta (125g/pers for a meal)

-Need more flour (just on pack more)

-Need more milk

-Less sugar

-Less tofu

-Soup are really good as starter

-Quantity of other things are perfect I think